OpenStudy (anonymous):

Given that the cost of a firm is C(q)=60q-12q^2+q^3. at what level of output is average cost minimum. and verify at the same level of output at which average cost is minimum average cost is equal to marginal cost

OpenStudy (dumbcow):

avg cost= C(q)/q marginal cost = C'(q) to minimize a function, set its derivative equal to 0 avg cost = A(q)= 60 -12q+q^2 A'(q) = 2q -12 = 0 --> q = 6